when 1 g of water at 100 deg. gets converted into steam at the same temp. thhe change in volume is____________cc


PLs explain how to do this!!!

Grade:9

1 Answers

Chetan Mandayam Nayakar
12 years ago

water volume= 1cc, steam volume=1.694 m3/kg*1g=1.694*10^-3*10^6=1694cc